# Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths

### v1

@@ Acknowledgements.
M.P. and I.D. acknowledge the support of the Canada Research Chair Program and the Natural Sciences and Engineering Research Council of Canada (NSERC, funding reference number RGPIN-2018-05425). This research has made use of NASA’s Astrophysics Data System Bibliographic Services. The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan as well as Princeton University. The HSC instrumentation and software were developed by the National Astronomical Observatory of Japan (NAOJ), the Kavli Institute for the Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo, the High Energy Accelerator Research Organization (KEK), the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan (ASIAA), and Princeton University. Funding was contributed by the FIRST program from the Japanese Cabinet Office, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), the Japan Society for the Promotion of Science (JSPS), Japan Science and Technology Agency (JST), the Toray Science Foundation, NAOJ, Kavli IPMU, KEK, ASIAA, and Princeton University. This paper makes use of software developed for the Large Synoptic Survey Telescope (LSST). We thank the LSST Project for making their code available as free software at http://dm.lsst.org. This paper is based [in part] on data collected at the Subaru Telescope and retrieved from the HSC data archive system, which is operated by Subaru Telescope and Astronomy Data Center (ADC) at National Astronomical Observatory of Japan. Data analysis was in part carried out with the cooperation of Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan.
@@ Appendix A Comparison with previous size estimates
Analysis of SourceXtractor++ performance guarantees the reliability and the essential absence of bias of our measurements (see Sect. 3.2, and Euclid Collaboration et al. 2023). Here we show a sanity check of our results by comparing our findings with independent measurements of morphological parameters derived for the same sample by Damjanov et al. (2023), that use SExtractor to fit a single-Sérsic model to the HSC/SSP $i-$band photometry of the HectoMAP sample.
Figure 13 shows the results for the HectoMAP mass-complete sample of quiescent galaxies with $R_{e}$ larger than half the median $i$-band seeing ($0.55$ arcsec). From left to right, the upper row shows the ratio $R_{e}/R_{e}^{D23}$ between our sizes $R_{e}$ and the corresponding sizes $R_{e}^{D23}$ by Damjanov et al. (2023) as a function of $M_{\star}$, $R_{e}^{D23}$, and $r_{petro}$, respectively. The middle and lower rows show the same for Sérsic index, $n/n^{D23}$, and axis ratio, $(b/a)/(b/a)^{D23}$, respectively. In each panel, black line and error bars show median and the interquartile ranges in ten equally-sampled bins of the corresponding $x$-axis.
@@
Figure 13: Comparison with previous measurements for HectoMAP galaxies in $i-$band. The upper row shows the ratio between our $R_{e}$ SourceXtractor++ measurements and the corresponding size $R_{e}^{D23}$ obtained by (Damjanov et al. 2023) using the SExtractor software. The middle and lower rows show the same, but for $n$ and $b/a$, respectively. From left to right, the panels show the dependence on $M_{\star}$, $R_{e}^{D23}$, and $r_{petro}$. Blue regions show the density distribution obtained with a gaussian kernel estimator. The darkest (smallest) region encloses the 10% of the sample, while the lightest (broadest) region includes the 90% of the sample; successive levels are separated by equal increases of $+10\%$ of the total sample. The black lines connect the medians of the ratios in ten equally-sampled bins of the corresponding $x$-axix. Error bars show the interquartile range of the corresponding bin.

### v2

@@ Acknowledgements.
We thank the anonymous referee for providing insightful comments and suggestions. M.P. and I.D. acknowledge the support of the Canada Research Chair Program and the Natural Sciences and Engineering Research Council of Canada (NSERC, funding reference number RGPIN-2018-05425). This research has made use of NASA’s Astrophysics Data System Bibliographic Services. The Hyper Suprime-Cam (HSC) collaboration includes the astronomical communities of Japan and Taiwan as well as Princeton University. The HSC instrumentation and software were developed by the National Astronomical Observatory of Japan (NAOJ), the Kavli Institute for the Physics and Mathematics of the Universe (Kavli IPMU), the University of Tokyo, the High Energy Accelerator Research Organization (KEK), the Academia Sinica Institute for Astronomy and Astrophysics in Taiwan (ASIAA), and Princeton University. Funding was contributed by the FIRST program from the Japanese Cabinet Office, the Ministry of Education, Culture, Sports, Science and Technology (MEXT), the Japan Society for the Promotion of Science (JSPS), Japan Science and Technology Agency (JST), the Toray Science Foundation, NAOJ, Kavli IPMU, KEK, ASIAA, and Princeton University. This paper makes use of software developed for the Large Synoptic Survey Telescope (LSST). We thank the LSST Project for making their code available as free software at http://dm.lsst.org. This paper is based [in part] on data collected at the Subaru Telescope and retrieved from the HSC data archive system, which is operated by Subaru Telescope and Astronomy Data Center (ADC) at National Astronomical Observatory of Japan. Data analysis was in part carried out with the cooperation of Center for Computational Astrophysics (CfCA), National Astronomical Observatory of Japan.
@@ Appendix A Comparison with previous size estimates
Analysis of SourceXtractor++ performance guarantees the reliability and the essential absence of bias of our measurements (see Sect. 3.2, and Euclid Collaboration et al. 2023). Here, we show a sanity check of our results by comparing our findings with independent measurements of morphological parameters derived for the same sample by Damjanov et al. (2023), that use SExtractor to fit a single-Sérsic model to the HSC/SSP $i$-band photometry of the HectoMAP sample.
Figure 13 shows the results for the HectoMAP mass-complete sample of quiescent galaxies with $R_{e}$ larger than half the median $i$-band seeing ($0.55$ arcsec). From left to right, the upper row shows the ratio $R_{e}/R_{e}^{D23}$ between our sizes $R_{e}$ and the corresponding sizes $R_{e}^{D23}$ by Damjanov et al. (2023) as a function of $M_{\star}$, $R_{e}^{D23}$, and $r_{petro}$, respectively. The middle and lower rows show the same for Sérsic index, $n/n^{D23}$, and axis ratio, $(b/a)/(b/a)^{D23}$, respectively. In each panel, black line and error bars show median and the interquartile ranges in ten equally sampled bins of the corresponding $x$-axis.
@@
Figure 13: Comparison with previous measurements for HectoMAP galaxies in $i$-band. The upper row shows the ratio between our $R_{e}$ SourceXtractor++ measurements and the corresponding size $R_{e}^{D23}$ obtained by (Damjanov et al. 2023) using the SExtractor software. The middle and lower rows show the same, but for $n$ and $b/a$, respectively. From left to right, the panels show the dependence on $M_{\star}$, $R_{e}^{D23}$, and $r_{petro}$. Blue regions show the density distribution obtained with a Gaussian kernel estimator. The darkest (smallest) region encloses the 10% of the sample, while the lightest (broadest) region includes the 90% of the sample; successive levels are separated by equal increases of $10\%$ of the total sample. The black lines connect the medians of the ratios in ten equally sampled bins of the corresponding $x$-axix. Error bars show the interquartile range of the corresponding bin.